OpenStudy (anonymous):

Factorize the following expression please. x to the power of 4-16

OpenStudy (lgbasallote):

\[x^4 - 16\] that's difference of two squares

OpenStudy (lgbasallote):

Hint: \[(a^2 - b^2) = (a+b)(a-b)\]

OpenStudy (anonymous):

x^4-16=(x^2+4)(x^2-4)=(x^2+4)(x+2)(x-2)

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow x^4 - 4^2 }\) \(\Large \color{MidnightBlue}{\Rightarrow (x^2)^2 - 4^2 }\) Try to factorize this by this identity: a^2 - b^2 = (a + b)(a - b)

OpenStudy (lgbasallote):

lol @ParthKohli

Parth (parthkohli):

What? @lgbasallote

OpenStudy (anonymous):

Thank you. I just was not sure about (x²+4). Have a nice day.

Parth (parthkohli):

Yw :D :P

OpenStudy (anonymous):

welcome